Boris Tobotras

                          boris@xtalk.msk.su

Summary

• Understanding of enterprise-wide information systems’ principles and implementation practices. Significant experience in system and application software design and development. Experience in local and wide area networks. Understanding of information systems security principles and implementation methods. Knowledge in tasks and solutions of information systems management. Experience in work in group. Able to be a team leader. Fluent English.

• Knowledge of theory and practical experience in development and implementation of structured document management systems. Understanding of SGML and XML technology, knowledge in SGML tools.

• Operating systems: RT-11, RSX-11, MS DOS, Windows, Interactive UNIX, AT&T UNIX SVR3, Solaris, AIX, Linux; detailed knowledge and practical experience in UNIX systems and TCP/IP networks administration, usage and programming.

• Platforms: DEC PDP-11, Besta-88, IBM PC, Sun SPARC, IBM RS/6000 and SP. Clustering systems. Disk arrays.

• High availability systems. Backup and archive software. Systems management systems.

• Programming language: Scheme and Lisp, Tcl/Tk, C, C++, Java, Unix shell scripting, PostScript, FORTRAN, Pascal, assemblers.

• Information security solutions. Firewalls: Checkpoint (Solstice) Firewall-1, TIS fwtk and Gauntlet, “Zastava-Jet”.

• Technologies: SGML/XML/HTML, DSSSL, HyTime, scripting, CGI, TeX, PDF.

Employment history

1994 — now : Jet Infosystems Position: service engineer, consultant.

Installation, service, customer support, maintenance, performance tuning, consulting and pre-sales in following areas:

• Full line of Sun workstations and servers: SPARCstation 5 to UltraEnterprise 6000;

• Storage systems: SPARCstorage Arrays, StorEdge, tape- and CD-libraries, RAID;

• OS Solaris and applications;

• clustering systems: SPARCcluster PDB; • high availability solutions: Qualix HA+, Veritas FirstWatch, Veritas Filesystem and

Volume Manager;

• backup and archive software: Legato (Solstice) NetWorker, Cheyenne (CA) ARCserve/Open;

• systems management systems: CA Unicenter TNG, SunNet Manager;

• “TurboSwift” SWIFT terminal.

Enterprise-wide projects on information security and information systems. Base platform: UNIX and Web.

Design, development and implementation of structured document management system, SGML-based. System is currently in active use in the company.

Intranet technology projects: design and development. CGI programming, fulltext search system with Informix DBMS access.

LAN, application and file servers’ administration.

1992-1994 : JSC “Uwertt”, JSC “RT”

Position: developer, software engineer.

Development of data compression module.

Development of portable (DOS/Windows/UNIX) library for interprocess communications.

Design and development of “intelligent editor”: document preparation system integrating editor, database for documents and objects used in documents, query facilities. System was developed up to sales ready state.

Custom module was integrated into NCSA telnet package (DOS TCP/IP stack), package was adapted for customer’s needs and documented.

1989-1992 : Moscow Telecommunication Institute, Research and Development facility, 15th laboratory Position: technician, engineer.

Development of software: digital filters’ modeling and characteristics calculations, special (polarization) processing of digital images, results visualization.

Non-standard peripherals (CCD-cameras, image accumulator, photo-printer) interfaces: adapters design and implementation (for LSI-11, IBM PC/ISA bus), device drivers programming.

Education

1987-1992 : Moscow Telecommunication Institute (currently Moscow Technical University of Communication and Computer Science) Automatics, telemechanics and electronics faculty (currently Radiotechnics).

Courses and trainings

Nov 7-11, 1994 : Sun Microsystems Computer Company, Moscow

Advanced SE Certification Course Nov 13-17, 1995 : SunService, Moscow

Solaris System Administration II

Nov 20-22, 1995 : SunU Course and Development, Moscow Sun Disk Management

Apr 29-30, 1996 : Sun Microsystems, Milpitas Volume Management with a SPARCstorage Array

May 1-3, 1996 : Sun Microsystems, Milpitas SPARCcluster Parallel Database Administration

Oct 21-23, 1996 : IBM, Moscow AIX Performance Management and Tuning

Nov 25-29, 1996 : IBM, Moscow AIX for UNIX Professionals

Jan 27-31, 1997 : IBM Central Europe & Russia Inc, Vienna SP Overview and Implementation Workshop

Apr 25-26, 1997 : Computer Associates International, Islandia Unicenter TNG Agents Programming

Feb 3-4, 1998 : Trusted Information Systems, Inc, Moscow Gauntlet Firewall: installation, configuration and management course

Personal information

27, married, two children. Non-smoker. Living in Moscow.